# Hierarchical Clustering Algorithms on Poisson and Cox Point Processes

### v2

@@ Abstract
Clustering is a widely used technique in unsupervised learning to identify groups within a dataset based on the similarities between its elements. This paper introduces three new hierarchical clustering models, Clustroid Hierarchical Nearest Neighbor ($\mathrm{CHN}^{2}$), Single Linkage Hierarchical Nearest Neighbor ($\mathrm{SHN}^{2}$), and Hausdorff (Complete Linkage) Hierarchical Nearest Neighbor ($\mathrm{H}^{2}\mathrm{N}^{2}$), all designed for datasets with a countably infinite number of points. These algorithms proceed through multiple levels of clustering and construct clusters by connecting nearest-neighbor points or clusters, but differ in the distance metrics they employ (clustroid, single linkage, or Hausdorff, respectively). Each method is first applied to the homogeneous Poisson point process on the Euclidean space, where it defines a phylogenetic forest, which is a factor of the point process and therefore unimodular. The results established for the $\mathrm{CHN}^{2}$ algorithm include the almost-sure finiteness of the clusters and bounds on the mean cluster size at each level of the algorithm. The mean size of the typical cluster is shown to be infinite. Moreover, the limiting structure of all three algorithms is examined as the number of levels tends to infinity, and properties such as the one-endedness of the limiting connected components are derived. In the specific case of $\mathrm{SHN}^{2}$ on the Poisson point process, the limiting graph is shown to be a subgraph of the Minimal Spanning Forest. The $\mathrm{CHN}^{2}$ algorithm is also extended beyond the Poisson setting, to certain stationary Cox point processes and and more generally to all stationary point process with no second-order descending chains. Similar finite-cluster properties are shown to hold in these cases. It is also shown that efficient detection of Cox-triggered aggregation can be achieved through this clustering algorithm.
@@ 1 Introduction
Clustering is a popular method in unsupervised learning, which plays an important role in data science. Given a representation of the objects, clustering aims at identifying groups, or clusters, based on some similarity measure between them. Usually, objects are represented as a finite number of points in $\mathbb{R}^{d}$.
Clustering algorithms can be organized into two broad categories: point assignment and hierarchical. In point assignment algorithms, some clusters are predefined, points are considered in some order, and each one is assigned to the cluster into which it best fits. Here, an important parameter is the number of clusters. The best known point assignment algorithm is the $k$-means algorithm (see [26] for a definition and further details). Hierarchical algorithms start with each point forming its own cluster. Clusters are then connected/merged on the basis of their similarities and form new clusters. The process of connecting and merging clusters is repeated until one reaches a single cluster to which all the points belong. A key component of hierarchical clustering is the function used to measure distances between clusters. This includes minimum distance (single linkage), Hausdorff distance (complete linkage), and centroid/clustroid distances. In centroid/clustroid distances, one or more representative points within each cluster, referred to as centroid or clustroid [^1] When the representative point of a cluster is not precisely located at the center of the cluster’s data, the term clustroid is used instead of centroid., are selected to collectively represent the entire cluster. The distance between clusters is then calculated on the basis of the distances between their respective centroids (clustroids). For more details, see [21] and [15].
The aim of the present paper is to study the properties of a class of natural clustering algorithms that essentially belong to the hierarchical category but also share features of point assignment. This class comprises the three hierarchical clustering algorithms listed in the abstract, namely the $\mathrm{CHN}^{2}$, $\mathrm{SHN}^{2}$, and $\mathrm{H}^{2}\mathrm{N}^{2}$ clustering algorithms. These algorithms follow the same overall principles for constructing clusters at each level, but differ in the distance used between clusters: $\mathrm{CHN}^{2}$ relies on clustroid distance, $\mathrm{SHN}^{2}$ uses single linkage distance, whereas $\mathrm{H}^{2}\mathrm{N}^{2}$ employs the Hausdorff (complete linkage) distance. The main focus of this paper is on the properties of $\mathrm{CHN}^{2}$. Some properties of $\mathrm{SHN}^{2}$ and $\mathrm{H}^{2}\mathrm{N}^{2}$ are however also discussed in Section 3.
In what follows, we give a high-level description of the clustering algorithms, centered on the $\mathrm{CHN}^{2}$ case, where the clustroid of a cluster is chosen as the pair of points that achieves the minimum distance in the cluster. More details on the definition of $\mathrm{CHN}^{2}$ are provided in Section 2. The variants associated with $\mathrm{SHN}^{2}$ and $\mathrm{H}^{2}\mathrm{N}^{2}$ are described in Section 3. Below, $D$ is the dataset, consisting of a countable collection of points in $\mathbb{R}^{d}$. The precise underlying assumptions will be spelled out later. They will for instance be satisfied by the collection of atoms of a homogeneous Poisson point process of the Euclidean space.
At level zero, each point in $D$ connects to its nearest neighbor point. Under these assumptions, this defines a collection of finite connected components called clusters of order 0 on the data set. Each cluster is made up of two directed trees, called cluster subtrees of order 0, that are connected to a directed cycle containing a pair of mutually nearest neighbor points (MNN). Call each such pair a cycle of order 0 and the points belonging to a cycle of order 0 cluster heads of order 0. See Figure 1  [^2] The graph constructed at this level is classical and called the nearest neighbor graph on $D$. Nearest neighbor graphs have been extensively studied in computational geometry ([12], and [13]) and play an important role in clustering and classification methods (see [9], [21], and [16]). They are known to capturing local proximity structures effectively, which makes them valuable in manifold learning, spatial data analysis, and other tasks where local neighborhoods are of interest (see [7])..
@@
At level one each 0-cycle is connected to its nearest 0-cycle. The pseudo-distance between two cycles is defined as the minimum (or infimum) distance between any point in the first 0-cycle and any point in the second. In the literature on clustering algorithms, this pseudo-distance is called the single linkage distance. The last name will be used in this paper to indicate this pseudo-distance. More precisely, define the exit point of a 0-cycle as the point in this 0-cycle that is closest to the nearest 0-cycle. For each exit point, replace the initial directed connection to the other point in its own cycle by a directed connection to the closest point in the nearest 0-cycle. This creates a directed graph on cluster heads of order 0, with finite connected components. Each component in this graph consists in directed trees whose nodes are cluster heads of order 0. Each tree is connected to a cycle containing a pair of 0-cluster heads that belong to two 0-cycles which are mutual nearest neighbors for single linkage. Each connected component of this graph, along with the cluster subtrees of order zero attached to them, defines a cluster of order 1 on the dataset. Each order 1 cluster contains some directed trees, called cluster subtrees of order 1. The nodes of these trees are either non-MNN points or cluster heads of order 0. Each such tree is connected to a cycle of length 2, which connects two cluster heads of order 0 which belong to two mutual nearest neighbor 0-cycles. This pair forms a cycle of order 1 and the points belonging to such a cycle will be called cluster heads of order 1. The latter are hence a subclass of that of cluster heads of order 0. See Figure 2.
Figure 2: Clusters of order one generated by the $\mathrm{CHN}^{2}$ clustering algorithm on a finite dataset. Each cycle of order zero is connected to its nearest neighbor pair with respect to the single linkage distance. The green arrows indicate the new connections of the vertices that are exit points of a cycle of order zero. The pink points represent the cluster heads of order 1, which are the exit points of the mutual nearest neighbor cycles with respect to the single linkage distance. The details of the elements within one cluster are shown in the figure.
By iterating this procedure, one can proceed hierarchically as above to define exit points of order 1 or higher, clusters of order 2 or higher, cluster subtrees of order 2 or higher, as well as cycles and cluster heads of order 2 or higher. For all $k\geq 0$, this construction results in a directed graph of order $k$ where each component or cluster of order $k$ is finite, made of a collection of cluster subtrees of order $k$, connected to a $k$-cycle connecting two cluster heads of order $k$. A cluster head of order $k$ is an exit point of order $k-1$ that connects two mutual nearest neighbor cycles of order $k-1$ with respect to the single linkage distance. Cluster heads of order $k$ hence form a subclass of those of order $k-1$.
All classification methods receive a phylogenetic interpretation, where nodes are seen as species. In the $\mathrm{CHN}^{2}$ clustering algorithm, the directed graph of order $k$ organizes species into an infinite collection of finite $k$-clusters. Each cluster contains directed trees (subtrees of order $k$) that specify parent-child relationships among species, up to cluster heads of order $k$, which can be considered the archetypes of the cluster. For these cluster heads, it is not possible to determine the parent species without increasing the order $k$.
Here are a few observations on $\mathrm{CHN}^{2}$ clustering:
  • It is defined as a hierarchical algorithm; however, it can also be considered a point assignment algorithm because the cluster heads of a given order are, in a sense, predefined through a mechanism that does not require the whole dataset but only the cycles of the previous order and the computation of the mutual nearest neighbor pairs of cycles.
  • For all $k$, the cluster heads (which can be seen as archetypes) of level $k$ are a deterministic functional of the data set. Their definition involves no extra randomness, as in, for example, the techniques based on the thinning of point processes described in more detail in Section 6.
  • $\mathrm{CHN}^{2}$ clustering is applicable to infinite data sets. Additionally, using the pseudo-distance between pairs results in limited computational effort, making it well-suited for large datasets. This feature distinguishes it from other known distance-based hierarchical clustering algorithms, such as the nearest-neighbor chain (see [8] and [18]).
  • Furthermore, the $\mathrm{CHN}^{2}$ clustering algorithm supports parallelization, allowing trees to be built in two sufficiently separated regions simultaneously. This enables faster processing and improved scalability.
As mentioned above, the aim of this paper goes beyond introducing new hierarchical clustering algorithms. It also studies how the algorithms behave when applied to point processes. Here is a summary of the main analytical results:
First, in Section 2, the $\mathrm{CHN}^{2}$ algorithm is considered on a homogeneous Poisson point process (PPP). It is shown that the algorithm can be described as a sequence of deterministic point-shifts on the PPP, with each point-shift representing a step of the algorithm (for more details on point-shifts, see [5]). At each step $k$, the point-shift constructs a random directed graph on the PPP. One key property is that the connected components (or clusters) of the directed graph of each level $k\in\mathbb{N}$ of the algorithm are all a.s. finite, as shown in Theorem 2.9. This theorem is proven in Section 5.1 by introducing the notion of Second-Order Descending Chain, which generalizes that of a descending chain in point processes (see [11] and [22]). In this chain, the distance between two consecutive points $x_{k}$ and $x_{k+1}$, denoted as $d_{k}$, satisfies the relation $d_{k}\leq\max(d_{k-1},d_{k-2})$, for all $k\geq 3$. It is shown that the PPP does not admit an infinite second-order descending chain a.s. This is shown to imply that every level of the $\mathrm{CHN}^{2}$ clustering algorithm contains only finite clusters. So the algorithm remains well-defined as it transitions from level $k$ to level $k+1$. Moreover, by considering the exit points at level $k$ as representatives of each cluster, it is shown that the intensity of the exit points at level $k$ is at most $\rho/2^{k+1}$, where $\rho$ is the intensity of the cluster heads of level 0 (which is known in closed form). Simulation results indicate that this upper bound can be further refined to approximately $\rho/3^{k+1}$ (see Figure 5). This result provides insight into the speed at which clusters merge at each step of the algorithm, forming new clusters at subsequent levels, see Remark 2.6.
Moreover, at each level, the resulting random graph is unimodular (see [2]). Corollary 2.8 further shows that, as $k\to\infty$, the limiting point-shift called Clustroid Hierarchical Nearest Neighbor $(\mathrm{CHN}^{2})$ Point-Shift exists. This limiting (point-shift) graph is called the Clustroid Hierarchical Nearest Neighbor Eternal Family Forest ($\mathrm{CHN}^{2}$ EFF). This infinite graph is unimodular under the Palm probability of $\Phi^{0}$.
The $\mathrm{CHN}^{2}$ EFF forms an infinite spanning forest or tree on the PPP. All connected components of $\mathrm{CHN}^{2}$ EFF have the same distribution (Proposition 2.11), and Corollary 2.12 shows that each connected component in $\mathrm{CHN}^{2}$ EFF is one-ended. These results hold for any stationary point process that does not admit a second-order descending chain almost surely (see Sections 2.2 and 5.1 for further details).
As mentioned above alternative definitions of the distance between clusters are also possible. Specifically, in Section 3, the two variants of the algorithm alluded to above are introduced more precisely: the Single Linkage Hierarchical Nearest Neighbor ($\mathrm{SHN}^{2}$) and the Hausdorff (Complete Linkage) Hierarchical Nearest Neighbor ($\mathrm{H}^{2}\mathrm{N}^{2}$) algorithms, which employ single linkage and Hausdorff distances, respectively. The properties of these algorithms on Poisson point processes can be analyzed by techniques similar to those of the $\mathrm{CHN}^{2}$ clustering algorithm.
In these cases, it is necessary to generalize the definition of the algorithm at each level to handle potentially infinite clusters. The random graphs constructed at each level of the $\mathrm{SHN}^{2}$ and $\mathrm{H}^{2}\mathrm{N}^{2}$ algorithms are unimodular, and their weak limits exist. These limits are referred to as the $\mathrm{SHN}^{2}$ Spanning Forest and the $\mathrm{H}^{2}\mathrm{N}^{2}$ Spanning Forest, respectively (see Theorem 3.2). To prove the existence of these forests, a sequence of point-shifts on the corresponding graphs is defined in Section 5.1. It is shown that the weak limits of these point-shifts exist and that the associated graphs are one-ended (see Remark 5.12). Moreover, Corollary 3.4 shows that, on the PPP, the $\mathrm{SHN}^{2}$ Spanning Forest is a subgraph of the Minimal Spanning Forest (MSF) introduced by Aldous and Steele in [23]. This result leads to a new class of open problems on the $\mathrm{SHN}^{2}$ Spanning Forest, potentially offering further insights into the behavior of the MSF on the PPP (see Section 7).
One of the generic goals of all clustering algorithms is to detect aggregation of points in a dataset. Working with a PPP is useful in this context because the PPP can be viewed as a null hypothesis, i.e., a baseline model that assumes no underlying aggregation other than those triggered by the fluctuations of the homogeneous PPP. To examine the behavior of the $\mathrm{CHN}^{2}$ algorithm in scenarios where aggregation does occur, Section 4 applies this algorithm to instances of Cox point processes, which exhibit noticeable aggregation (see Figures 11 and 12). It is shown that, in these cases, the $\mathrm{CHN}^{2}$ algorithm enjoys properties analogous to those observed under the PPP, such as the finiteness of clusters at each hierarchical level and the weak convergence of the limiting graph. Furthermore, it is shown how the $\mathrm{CHN}^{2}$ clustering algorithm effectively detects aggregation in the dataset.
@@ 2 Clustroid Hierarchical Nearest Neighbor Clustering on a Point Process
This section considers the $\mathrm{CHN}^{2}$ algorithm on the homogeneous PPP. As previously stated in the introduction, each level of the algorithm can be characterized as a point-shift on the PPP. A point-shift $f$ maps, in a translation invariant way, each point of a stationary point process $\phi$ to some point of $\phi$ (see [20], [24], and [4] for more details).
@@ 2.1 Construction of the Pre-limit Point-Shifts
$\delta(S,T)=\min_{x\in S,y\in T}d(x,y),$ (1)
where $d$ denotes the Euclidean distance. As mentioned in the introduction, this pseudo-distance is called the single linkage distance in the literature. This function is non-negative and symmetric. However, it does not satisfy the triangle inequality. Additionally, different sets have a pseudo-distance of 0 when they share at least one point.
@@
The first point-shift, $f^{0}$, is the nearest neighbor point-shift on $\Phi^{0}$ w.r.t. Euclidean distance $d$ between points. That is for all $x$, $f^{0}(x)={\mathrm{NN0}}(x)$, where NN0 maps a point to its nearest neighbor in $\Phi^{0}$.
@@
The graph of the point-shift $f^{0}$ (the $f^{0}$-graph) has all its connected components in the $\mathcal{F/F}$ class in the sense of the point-shift classification theorem in [6]. That is, all of its connected components are a.s. finite and have a unique cycle.
@@
Note that any infinite path in the $f^{0}$-graph on the PPP is a descending chain. A descending chain in a point process is an infinite sequence of points in the process where the distances between consecutive points form a decreasing sequence (see [22], [17], and [11]). It is shown in [11] that the homogeneous PPP does not admit any descending chains a.s. Therefore, the connected components of the $f^{0}$-graph are almost surely finite and belong to the $\mathcal{F/F}$ class of the point-shift classification theorem. Each connected component consists of directed trees connected to a single cycle that connects two mutual nearest neighbor (MNN) points of $\Phi^{0}$. ∎
@@
Each connected component of the $f^{0}$-graph on the PPP has a cycle of length two. These cycles are called 00-cycles. The connected components of the $f^{0}$-graph are also referred to as clusters of order zero. The vertices in the $f^{0}$-graph that belong to 00-cycles are called cluster heads of order zero. The union of cluster heads of order zero forms a sub-PP of $\Phi^{0}$ denoted by $\Phi^{0}_{c}$. Let $\Phi^{0}_{a}=\Phi^{0}\setminus\Phi^{0}_{c}$.
If the edges of the 00-cycles in the $f^{0}$-graph are deleted, one obtains a collection of directed trees referred to as the cluster subtrees of order 0. Each vertex $x$ in a cluster subtree of order 0 is either a cluster head of order zero (i.e., $x\in\Phi^{0}_{c}$) or is such that $x\in\Phi^{0}_{a}$, and there is a path from $x$ to a cluster head of order 0. See Figure 3.
Figure 3: Left picture: The $f^{0}$-graph generated on PPP in blue with the 00-cycles shown in bold;
Right picture: Cluster subtrees of order 0 obtained by deleting the edges of the 00-cycles (clusters consisting of a single point are not visible here). In each cluster subtree, one vertex is a cluster head of order zero, and the other vertices have directed paths towards the cluster head of order zero within the cluster.
@@
Let $\{C^{0}_{i}\}_{i}$ denote the collection of 00-cycles, and let $\{S^{0}_{i}\}_{i}$ be the collection of pairs of points in these cycles, namely $S^{0}_{i}$ is the pair of head points belonging to the cycle $C^{0}_{i}$. The sets $\{S^{0}_{i}\}_{i}$ form a translation-invariant partition of the support of $\Phi^{0}_{c}$.
@@
[…]
@@
[…]
@@
[…]
@@
  • The ${k-1}$-cycles $\{C^{k-1}_{i}\}$ of the connected components are all of length 2 a.s.;

### v3

@@ Abstract
This paper introduces a hierarchical clustering algorithm, the Clustroid Hierarchical Nearest Neighbor ($\mathrm{CHN}^{2}$), designed for datasets with a countably infinite number of points. The method builds clusters across successive levels by linking nearest-neighbor points or clusters using the clustroid distance. The properties of this algorithm make it suitable for very large datasets.
To evaluate its properties, we first apply the algorithm to the homogeneous Poisson point process, which serves as a natural null-hypothesis model with no intrinsic aggregation. In this setting, the algorithm generates a random forest that is a factor of the Poisson point process and hence unimodular. We prove that at every level, the level-$k$ graph has only finite connected components (a.s.) and derive bounds on their mean size. We also establish the existence of a limiting graph as the number of levels tends to infinity. In this limit, clusters are infinite and one-ended, which induces a natural order within each component and supports a tree-like phylogenetic interpretation.
Beyond the Poisson case, we extend the analysis to a class of Cox and more general stationary point processes without second-order descending chains (introduced here), for which analogous results hold. Simulations show that comparing these cases with the Poisson baseline allows an efficient detection of aggregation, thereby linking the stochastic-geometric analysis to practical clustering tasks.
keywords:
Hierarchical Clustering , Clustering on Poisson Point Process , Cox Point Process , Random Trees , Unimodular Random Graphs
[…]
[1]organization=INRIA/ENS Paris
@@ 1 Introduction
Clustering is a core task in unsupervised learning: given a representation of objects, one seeks a structure formed by a similarity notion. Classical methods often target moderate-sized finite datasets. In contrast, this work introduces and analyzes a hierarchical distance–based algorithm that applies equally well to very large data sets and even to countably infinite point sets.
We begin with a brief reminder of hierarchical clustering. One starts from singleton components, specifies a notion of proximity between components, and repeatedly links the closest ones to form larger components. The outcome is a multiscale summary of the data. Common choices for between-component proximity include minimum interpoint distance (single linkage), maximum/Hausdorff distance (complete linkage), and centroid- or clustroid-based variants; see [Murtagh, Murtagh2]. The key modeling choice is how components are represented when they interact.
Our focus is on the following algorithm: the Clustroid Hierarchical Nearest Neighbor algorithm, denoted by $\mathrm{CHN}^{2}$. The name reflects two design decisions. First, each component is represented by a clustroid, i.e., a representative chosen from the data rather than an external centroid. Second, components are linked using a nearest-neighbor rule (the superscript “22” is shorthand for “NN,” so we write $\mathrm{CHN}^{2}$ instead of $\mathrm{CHNN}$). Informally, the algorithm works as follows. At level 0, every point connects to its nearest neighbor. This produces finite components organized around mutual nearest-neighbor (MNN) pairs; the two points in such a pair form the component’s clustroid pair. Under mild general-position assumptions (no distance ties) this pair is a.s. unique. At higher levels, each current component is summarized solely by its clustroid pair, and components are linked to their nearest other component with respect to the clustroid distance (the distance between clustroid pairs). The connected components of the resulting graph are the clusters at that level. Iterating this rule builds a hierarchy. Figure 1 illustrates the level-0 picture on a finite dataset; formal definitions are deferred to Section 2.
@@
A simple phylogeny picture is useful here. One can view clustroid pairs as archetypes of their components. Linking archetypes across levels organizes the data into a tree-like summary: fine distinctions appear near level 0, while broader relations appear at higher levels. This finite-data interpretation guides the intuition; a rigorous asymptotic counterpart will appear later when we study the Poisson model.
The design of the algorithm scales well to large data. Once a component forms, only its two clustroid points are needed to decide future links. There is no extra optimization step to find representatives—the clustroid pair arises automatically as an MNN pair inside the component. This autonomy reduces computation and memory, since decisions are made on representatives rather than on all points, and it also enables parallelization: widely separated regions can be processed independently and hence in parallel.
We now fix the terminology. In what follows, clusters are the connected components produced by the algorithm at a given level, whereas aggregation refers to distinguishable group structure in the data itself (in the sense used in data analysis).
To study properties at the scale of very large or infinite datasets, we place the algorithm in a point-process framework. As a baseline, we adopt a null model with no intrinsic aggregation and apply $\mathrm{CHN}^{2}$ to the homogeneous Poisson point process (PPP). It will be shown that, in this setting, the algorithm produces a random forest that is a factor of the Poisson point process and hence unimodular (so the mass-transport arguments apply; see [Aldous] for more details). Unimodularity is important because it turns global properties into local expectations and gives structural leverage: for instance, in a unimodular random graph, the Foil Classification Theorem (see page 10 in [Baccelli2017]) classifies the size and the number of ends in the clusters. In our case, we prove that, at every level, all clusters are almost surely finite; we derive bounds on their mean sizes. We also establish the existence of a limiting graph as the number of levels tends to infinity. we prove that the mean size of the typical cluster in this limiting graph is infinite and all the clusters are one-ended. Returning to the phylogeny picture, one-endedness yields a natural order within the data (species) in each infinite cluster.
It is worth emphasizing that these finiteness and structural results are far from trivial in the infinite setting. In finite datasets, clusters produced by hierarchical algorithms are necessarily finite, but when the data form a stationary point process, the situation is subtler. Even the nearest-neighbor graph (the first level of the hierarchy) may contain infinite components; this occurs, for example, in certain non-Poisson stationary processes that admit infinite descending chains (see [Daley]). The proof that the homogeneous Poisson process yields only finite clusters thus requires specific geometric arguments and plays a key role in establishing the well-definedness of the hierarchy at all levels. Likewise, in the limiting infinite graph, the fact that all connected components are one-ended is not automatic: graphs with multiple ends can exhibit branching structures where no consistent order can be defined between distant points. The one-endedness property therefore ensures that, even in the infinite-data regime, each cluster admits a natural hierarchical order—providing the basis for its phylogenetic interpretation.
The main new methodological tool is the notion second-order descending chain introduced here. We show that the PPP admits almost surely no infinite second-order descending chain. This is instrumental in proving the finiteness of the level-$k$ clusters and yields intensity bounds for the level-$k$ representatives.
The approach is extended beyond the Poisson case. Cox point processes serve as prototypical models with aggregation. We introduce a class of Cox point processes for which all results established in the Poisson case hold. More generally, the results extend to all stationary point processes without second-order descending chains. In these settings, $\mathrm{CHN}^{2}$ has the same finiteness properties as the Poisson case. Simulations on the Cox processes show that, relative to the Poisson baseline, the algorithm efficiently detects aggregation, thereby distinguishing structural organization from purely stochastic dispersion.
Our contributions are the following: (i) the definition and study $\mathrm{CHN}^{2}$, a hierarchical nearest-neighbor algorithm that compresses components to clustroid pairs and scales to very large and countably infinite datasets; (ii) on the PPP, the proof of almost-sure finiteness at every level, bounds on mean cluster sizes, unimodularity of the level graphs, and the existence of a unimodular one-ended limiting graph; (iii) the introduction of the second-order descending chain method (iv) intensity bounds for representatives; and (v) the extention of the analysis to a class Cox and other stationary point processes without second-order descending chains (vi) the evidence for aggregation detection via simulations relative to the Poisson baseline.
The paper is organized as follows. Section 2 introduces the $\mathrm{CHN}^{2}$ algorithm on the homogeneous Poisson point process, defining its level-by-level construction through a sequence of point-shifts and the corresponding hierarchical clusters. This section also establishes the scale invariance of the construction, the finiteness property of each pre-limit graph, and the existence and unimodularity of the limiting graph, referred to as the $\mathrm{CHN}^{2}$ Eternal Family Forest (EFF). Section 3 extends the $\mathrm{CHN}^{2}$ clustering framework to Cox and other stationary point processes, illustrating its use for detecting point aggregation through numerical simulations and proposing stopping criteria for identifying hierarchical levels of aggregation. Section 4 provides the proofs of finiteness for all pre-limit point-shift graphs, showing that they contain no infinite paths by excluding the existence of second-order descending chains. Finally, Section 5 discusses several open questions and conjectures related to the connectivity and ergodic properties of the limiting graph, as well as the possible existence of a point-map probability for the $\mathrm{CHN}^{2}$ point-shift.
@@ 2 Clustroid Hierarchical Nearest Neighbor Clustering on a Point Process
This section considers the $\mathrm{CHN}^{2}$ algorithm on the homogeneous PPP. As previously stated in the introduction, each level of the algorithm can be characterized as a point-shift on the PPP. A point-shift $f$ maps, in a translation invariant way, each point of a stationary point process $\phi$ to some point of $\phi$ (see [Mecke], [Tho00], and [Pointshift] for more details).
@@ 2.1 Construction of the Pre-limit Point-Shifts
$\delta(S,T)=\min_{x\in S,\,y\in T}\mathfrak{d}(x,y),$ (1)
where $\mathfrak{d}$ denotes the Euclidean distance. As mentioned in the introduction, this pseudo-distance is called the single linkage distance in the literature. This function is non-negative and symmetric, but it does not satisfy the triangle inequality. Additionally, different sets have a pseudo-distance of zero whenever they share at least one point.
Preliminaries on unimodularity.
Before proceeding, let us recall that a unimodular graph is a random rooted graph $[G,o]$ that satisfies the Mass Transport Principle (MTP):
$\mathbb{E}\Big[\sum_{x\in V(G)}f(G,o,x)\Big]=\mathbb{E}\Big[\sum_{x\in V(G)}f(G,x,o)\Big],$
for all nonnegative measurable functions $f$ defined on doubly rooted graphs (see [Aldous, Baccelli2017]). Unimodularity ensures that global statements can be derived from local information. As mentioned in the introduction, this property is repeatedly used in our proofs. It allows us to apply general results such as the foil classification theorem for unimodular networks, and the NoInfinite/Finite Inclusion Theorem (see [Baccelli2017, Aldous] for details).
Foils. Given a point-shift $f$ on a (random) rooted graph $[G,o]$, define the equivalence relation $x\sim y\iff\exists\,m,n\geq 0\text{ such that }f^{m}(x)=f^{n}(y)$. The equivalence classes for $\sim$ are called foils (they can be viewed as level sets transverse to the flow induced by the parent map $f$). The Foil Classification Theorem [Baccelli2017] classifies components according to the finiteness/infiniteness of the component and of its foils (e.g., $\mathcal{F/F}$, $\mathcal{I/I}$).
@@
The first point-shift, $f^{0}$, is the nearest-neighbor point-shift on $\Phi^{0}$ with respect to the Euclidean distance $\mathfrak{d}$ between points. That is, for all $x\in\Phi^{0}$, we define $f^{0}(x)=\mathrm{NN0}(x)$, where $\mathrm{NN0}$ maps a point to its nearest neighbor in $\Phi^{0}$.
@@
The graph of the point-shift $f^{0}$ (the $f^{0}$-graph) is unimodular, and all its connected components belong to the $\mathcal{F/F}$ class in the sense of [Baccelli2017]; that is, each component is almost surely finite and contains exactly one cycle. Equivalently, both the component and its foils (equivalence classes induced by the point-shift) are a.s. finite.
@@
The unimodularity of the $f^{0}$-graph follows from the fact that it is obtained as a covariant shift on a stationary point process under the Palm probability (see [Baccelli2017, Aldous]). To see that all the connected components of this graph is finite, note that any infinite path in the $f^{0}$-graph on the PPP would form a descending chain, i.e., an infinite sequence of points whose consecutive distances form a decreasing sequence (see [Meester, Matching, Daley]). However, it is shown in [Daley] that the homogeneous PPP does not admit any descending chain almost surely. Therefore, the connected components of the $f^{0}$-graph are almost surely finite and consist of directed trees attached to a single cycle connecting two mutual nearest neighbors of $\Phi^{0}$. ∎
@@
Each connected component of the $f^{0}$-graph on the PPP has a cycle of length two. These cycles are called 0-cycles. The connected components of the $f^{0}$-graph are also referred to as clusters of order zero. The vertices in the $f^{0}$-graph that belong to 0-cycles are called cluster heads of order zero. The union of cluster heads of order zero forms a sub-PP of $\Phi^{0}$ denoted by $\Phi^{0}_{c}$. Let $\Phi^{0}_{a}=\Phi^{0}\setminus\Phi^{0}_{c}$.
If the edges of the 0-cycles in the $f^{0}$-graph are deleted, one obtains a collection of directed trees referred to as the cluster subtrees of order 0. Each vertex $x$ in a cluster subtree of order 0 is either a cluster head of order zero (i.e., $x\in\Phi^{0}_{c}$) or is such that $x\in\Phi^{0}_{a}$, and there is a path from $x$ to a cluster head of order 0. See Figure 2.
Figure 2: Left picture: The $f^{0}$-graph generated on PPP in blue with the 0-cycles shown in bold;
Right picture: Cluster subtrees of order 0 obtained by deleting the edges of the 0-cycles (clusters consisting of a single point are not visible here). In each cluster subtree, one vertex is a cluster head of order zero, and the other vertices have directed paths towards the cluster head of order zero within the cluster.
@@
Let $\{C^{0}_{i}\}_{i}$ denote the collection of 0-cycles, and let $\{S^{0}_{i}\}_{i}$ be the collection of pairs of points in these cycles, namely $S^{0}_{i}$ is the pair of head points belonging to the cycle $C^{0}_{i}$. The sets $\{S^{0}_{i}\}_{i}$ form a translation-invariant partition of the support of $\Phi^{0}_{c}$.
@@
The NN1 graph is a directed graph on $\{S^{0}_{i}\}_{i}$. There is an edge from $S^{0}_{i}$ to $\mathrm{NN1}(S^{0}_{i})$, the nearest neighbor of $S^{0}_{i}$ based on the single linkage distance. Let $\mathrm{NN1}(S^{0}_{i})=S^{0}_{j}$. There is a point, denoted $x_{i}$, in $S^{0}_{i}$ and a point, $x_{j}$, in $S^{0}_{j}$ that achieve the minimum in $\delta(S^{0}_{i},S^{0}_{j})$, denoted by $nn1(x_{i})=x_{j}$. The point $x_{i}$ is called the exit point of order zero in the pair $S^{0}_{i}$.
Define $f^{1}$ to be the point-shift on $\Phi^{0}$ that coincides with $f^{0}$ everywhere except for the exit points of order 0. For each such point, say $x\in S^{0}$, where $S^{0}$ is the pair where $x$ is the exit point, define $f^{1}(x)=\mathrm{nn1}(x)$. In other words, for each exit point $x$ of an $f^{0}$-cycle, $f^{1}$ is obtained from $f^{0}$ by replacing the initial image (which was the mutual nearest neighbor of $x$ with respect to $d$) by $\mathrm{nn1}(x)$.
@@
Each connected component of the graph of $f^{1}$ on $\Phi^{0}$ is referred to as a cluster of order 1. It will be shown in Corollary 4.6 that these connected components are almost surely $\mathcal{F/F}$, meaning that each cluster in the graph of $f^{1}$ is finite and contains a unique cycle. These cycles are of length 2 a.s. and connect two exit points of order 0, denoted $x$ and $y$, such that $\mathrm{nn1}(x)=y$ and $\mathrm{nn1}(y)=x$. These cycles are called $1$-cycles.
@@
The NN2 graph is a directed graph on $\{S^{1}_{i}\}_{i}$. There is an edge from $S^{1}_{i}$ to $\mathrm{NN2}(S^{1}_{i})$, the nearest neighbor of $S^{1}_{i}$ based on the single linkage distance. Let $\mathrm{NN2}(S^{1}_{i})=S^{1}_{j}$. There is a point, denoted $x_{i}$, in $S^{1}_{i}$ and a point, $x_{j}$, in $S^{1}_{j}$ that achieve the minimum in $\delta(S^{1}_{i},S^{1}_{j})$, denoted by $\mathrm{nn2}(x_{i})=x_{j}$. The point $x_{i}$ is called the exit point of order 1 in the pair $S^{1}_{i}$.
@@
  • The k1{k-1}-cycles $\{C^{k-1}_{i}\}$ of the connected components are all of length 2 a.s.;